OpenStudy (anonymous):

Which of the following represents the equation: 2x= x + 8? Two times x is equal to the product of x and 8. The sum of twice x and eight equals x. Twice the number x is eight more than x. X squared is the sum of x and 8.

OpenStudy (anonymous):

Twice the number x is eight more than x.

Parth (parthkohli):

@vinrar Nicely explained :)

OpenStudy (theviper):

yup

Parth (parthkohli):

And nicely spelled

OpenStudy (anonymous):

thank uu

Parth (parthkohli):

Just FYI - you are not allowed to post answers directly on this site. Okay?

Parth (parthkohli):

Please check this out. http://openstudy.com/code-of-conduct

OpenStudy (theviper):

2 times x= 2x & is eight more than x= x+8 2x=x+8

OpenStudy (anonymous):

Oh, im sorry..i was answering questions till now. ill b careful nxt onwards.